Selected amounts from Reznor Company’s trial balance as of 12/31/16 appear below: (Note that all of these accounts have their standard or normal debit or credit balance.)  No adjusting entries have been made yet this year.

Account

Balance

Accounts Payable

$360,000

Accounts Receivable

$346,000

Accumulated Depreciation- Equipment

$210,000

Allowance for Doubtful Accounts

$35,000

Cash

???

Common Stock

$66,000

Dividends Declared

$20,000

Dividends Payable

$10,000

Equipment

$900,000

Insurance Expense

$24,000

Interest Expense

$30,000

Merchandise Inventory

$300,000

Notes Payable (due 6/1/17)

$400,000

Prepaid Rent

$102,000

Retained Earnings

$412,000

Salaries and Wage Expense

$182,000

Subscription Revenue

$445,000

Supplies

$22,000

  • Prepare adjusting journal entries at year end, December 31, 2016, based on the following supplemental information.
    1. The equipment has a useful life of 8 years with a $60,000 salvage value. (Straight-line method being used).

2. The Note Payable is an 6% per year note that pays interest on the first day of January, April, July, and October.

3. The company took out insurance on its facilities for the first time on September 1st, paying for a year’s worth of insurance with a debit to a temporary account.

4. The rent payment of $102,000 covered the six months from November 30, 2016 through May 31, 2017.

5. The next payroll will be $4,000 and will be paid on January 7th for the two weeks ending January 7th.

6. Inventory of the supply room performed on 12/31/16 reported $6,000 in supplies on-hand.

  • If this company uses reversing entries, which of the above adjusting entries would be eligible to be reversed?  

Please show all of the work
Answer #1
Account Titles Debit Credit
Depreciation Expense $      105,000 =(900000-60000)/8
       Accumulated Depreciation $      105,000
Interest Expense $          6,000 =400000*6%*3/12
       Interest Payable $          6,000
Prepaid Insurance $        16,000 =24000/12*8
       Insurance Expense $        16,000
Rent Expense $        17,000 =102000/6*1
       Prepaid Rent $        17,000
Salaries and Wage Expense $          2,000 =4000/2
      Salaries and Wage Payable $          2,000
Supplies Expense $        16,000 =22000-6000
       Supplies $        16,000

Interest Paybale and Salaries and Wage Payable transaction requires reversing entries.

If you have any query, kindly comment with your query and please mark thumbs up.